OpenStudy (anonymous):

How many 4 inch by 8 inch bricks are needed to build a sidewalk 6 feet wide and 24 feet long?

OpenStudy (anonymous):

How would this be solved in the most simple method?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first convert the feet to inches for the sidewalk then figure out the area by multiplication

OpenStudy (misty1212):

the area of the brick is easy enough, it is \(8\times 4=32\) square inches

OpenStudy (ibrafacts):

yup u need to convert them

OpenStudy (misty1212):

once you find the area of the sidewalk in square inches, divide the result by \(32\)

OpenStudy (anonymous):

You have to convert 6ft and 24ft into inches so 6*12=72 and 24*12=288 and divide 72 by 6 and 288 by 24.